# ChartAnchor: Chart Grounding with Structural-Semantic Fidelity and Data Recovery

### v1

@@ A.2. Detailed Code Generation and Augmentation for Existing Datasets
A.3. Filtering Strategy and Statistics
@@
Table 7. Sample counts at each filtering stage.
@@
categorical4Pariscategorical4Aquantitative489.5
@@
Table 8. Review rubric for chart–table–code triples.
@@ Appendix B More Analysis about Dataset
Appendix C Broader Impacts
Appendix D Model License
@@
Table 9. Summary of licenses in models that are evaluated in ChartAnchor. Entries marked with “Not Applicable” indicate that authors do not have an explicit code license displayed within the codebase or model checkpoint page.
@@ Appendix E Experiments
E.1. Prompts
Figure 6. Prompt for Chart-to-Code task. The orange text in curly braces denotes a selectable visualization library.
Figure 7. Prompt for Controlled Chart-to-Table task. The orange text represents placeholders, which varies according to different table headers.
@@ E.2. Additional Analysis
- Color Feature Presents Unique Challenges in Visual Decoding. Color accuracy remains the lowest-scoring aspect within the Visual Structure Consistency metrics, even for top-performing models such as GPT-4o (34.64), Claude-3-7-Sonnet(40.76) and InternVL3-14B (26.13). This suggests that fine-grained color differentiation poses unique challenges for current visual encoders, especially in complex or low-contrast chart regions. In many cases, the visual abstraction processes used by these models may reduce sensitivity to precise pixel-level color features.
- Claude-3-7-Sonnet achieves the highest CLIP Score despite lower structural fidelity. Claude-3-7-Sonnet obtains the highest Clip Score (94.11), surpassing GPT-4o (86.88), indicating strong perceptual alignment with original chart styles. However, its overall Chart-to-Code accuracy (61.18) and structural metrics (e.g., Color: 40.79) remain lower, revealing a gap between visual realism and semantic correctness. This suggests that high CLIP similarity does not imply accurate structural grounding.
@@
- High Pass Rate but Low Fidelity: Another failure mode is models generating syntactically valid outputs that “mask critical failures” in content. For example, DeepSeek-VL-7B successfully produces parsable code nearly every time ( about 98% pass rate), yet its data extraction is almost nonexistent (strict F1 ¡1%). It often writes basic chart code that runs but does not capture the actual data or visual details. This strategy yields a high functional score but extremely low data fidelity, indicating the model is defaulting to trivial or placeholder outputs to avoid errors rather than truly understanding the chart.
@@
- A high CLIP score does not guarantee accurate data reconstruction. A model may generate a chart that appears visually correct—with the right chart type, structure, and even similar color distribution—while the underlying data is incorrect. For example, Claude-3-7-Sonnet achieves a high CLIP score (indicating strong visual similarity), but its data fidelity (e.g., F1) is only comparable to GPT-4 and slightly below Qwen2.5-VL-32B, which has the best high-tolerance F1 (39.6) but a much lower CLIP score (64.04). This shows that visual similarity and data accuracy can diverge. As noted in the introduction, a model may “reproduce a chart’s appearance while silently altering the data.” Thus, CLIP and data F1 are complementary: the former captures visual errors, the latter semantic ones. Both are essential for evaluating chart fidelity.
E.3. Analysis Across Chart Types
@@
Figure 8. Comparison of Overall Across Chart Types for Different Models on Chart-to-Code Task
@@
Figure 9. Comparison of Slight F1-Score Across Chart Types for Different Models on Controlled Chart-to-Table Task
@@ Appendix F Case Study of ChartAnchor
Figure 10. Image Example1 for Chart-to-Code Task.
@@
Figure 11. Image Example2 for Chart-to-Code Task.
@@
Figure 12. Image Example3 for Chart-to-Code Task. The gray image indicates that the model did not generate a valid image or the code parsing failed
@@
Figure 13. Image Example4 for Chart-to-Code Task. The gray image indicates that the model did not generate a valid image or the code parsing failed
@@
Figure 14. Image Example5 for Chart-to-Code Task. The gray image indicates that the model did not generate a valid image or the code parsing failed
@@
Figure 15. Image Example6 for Chart-to-Code Task.
@@
Figure 16. Image Example7 for Chart-to-Code Task. Each subfigure shows the output from a different model.
@@
Figure 17. Image Example8 for Chart-to-Code Task. The gray image indicates that the model did not generate a valid image or the code parsing failed.
@@
Example 9: From the gold table (Table 10) and the generated tables (Figure 18), by comparing the row structures and numeric alignments, we observe that claude-3-7-Sonnet produces a table closely aligned with the gold table, with all values either correct or slightly off. Qwen2.5-VL-7B shows noticeable deviations in the ”RDS 18–49” and ”RDS Total” columns for 2014 and 2016. Qwen2.5-VL-32B gets closer but still introduces uniform values in 2015 that deviate from the gold data. Claude-3-7-Sonnet demonstrates the highest data fidelity and variation consistency across rows.
@@
Table 10. Example 9: Gold Table
Figure 18. Example9. Gold Image and Generated Tables for Controlled Chart-to-Table Task.
@@
Table 11. Example 10: Gold Table
Example 10: From the gold table (Table 11) and the generated tables (Figure 19), we can see, in this example, Claude-3-7-Sonnet is the only model that correctly identifies the categorical theta labels (e.g., AL29, GD35) and extracts reasonable numerical values for both ”Today-r” and ”Mean-r,” achieving the highest alignment with the gold table. In contrast, Qwen2.5-VL-7B and Qwen2.5-VL-32B fail to capture the label names and instead generate numerical theta angles, indicating a misunderstanding of the scatterpolar chart structure. Their values also significantly diverge from the reference data, limiting their usability for structured analysis.
Figure 19. Example10. Gold Image and Generated Tables for Controlled Chart-to-Table Task.
Example 11: From the gold table (Table 12) and the generated tables (Figure 20), we can see, Claude-3-7-Sonnet is the only model that accurately captures both the structure and values of the gold table. It correctly identifies incremental changes (including positive and negative values) and matches the final ”Price current” result. In contrast, Qwen2.5-VL-7B and Qwen2.5-VL-32B misinterpret the chart as cumulative values rather than stepwise deltas. As a result, their tables lose the core logic of a waterfall breakdown and deviate significantly from the reference.
@@
Table 12. Example 11: Gold Table
Figure 20. Example11. Gold Image and Generated Tables for Controlled Chart-to-Table Task.
@@ Appendix G Chart Examples
This section presents visual examples of the chart categories included in our benchmark. Each image illustrates a distinct chart type, showcasing the diversity in structure, data encoding, and visual design. Figures 21 display representative samples from all 30 categories.
Figure 21. Representative chart images in our benchmark.

### v3

@@ A.2 Detailed Code Generation and Augmentation for Existing Datasets
A.3 Filtering Strategy and Statistics
@@
Table 8: Sample counts at each filtering stage.
@@
categorical4Pariscategorical4A-
quantitative489.5
@@
Table 9: Review rubric for chart–table–code triples.
@@ Appendix B More Analysis about Dataset
Appendix C Model License
@@
Claude-4-5-Sonnet Proprietary Proprietary
Gemini-3-Pro Proprietary Proprietary
@@
Qwen3-VL-32B-Instruct Apache 2.0 Apache 2.0
GLM-4.5V-106B MIT Apache 2.0
Qwen3-VL-235B-A22B-Instruct Apache 2.0 Apache 2.0
Table 10: Summary of licenses in models that are evaluated in ChartAnchor. Entries marked with “Not Applicable” indicate that authors do not have an explicit code license displayed within the codebase or model checkpoint page.
Appendix D Broader Impacts
@@ Appendix E Experiments
Figure 6: Prompt for Chart-to-Code task. The orange text in curly braces denotes a selectable visualization library.
E.1 Implementation Details
For all models, we set the temperature to $\tau=0.1$ and use top-$p$ sampling with $p=0.95$ for decoding. The maximum generation length is capped at 16384 tokens. For open-weight models, we adopt bfloat16 precision during inference. All experiments are conducted on H100 80G GPUs. Results are reported from a single run per setting, with low-temperature decoding ensuring stability.
E.2 Prompts
Figure 7: Prompt for Controlled Chart-to-Table task. The orange text represents placeholders, which varies according to different table headers.
@@ E.3 Additional Analysis
- Chart-to-Code Is Fundamentally Constrained by Long-Horizon Execution Requirements. Both DeepSeek-VL and DeepSeek-VL2 exhibit consistently low pass rates on the Chart-to-Code task (32.05/7.98), both operate under a maximum sequence length of 4096 tokens. Chart-to-Code requires long-horizon contextual coherence and strict execution fidelity, as successful generation depends on preserving object-level consistency across multiple code segments, including figure initialization, axis construction, and API-level correctness. Under limited context budgets, even minor inconsistencies in earlier object definitions can propagate into non-executable outputs. In contrast, tasks that do not enforce execution validity are far less sensitive to such long-context constraints. This indicates that Controlled Chart-to-Code inherently demands stronger long-context modeling and larger effective context windows, making it substantially more challenging than perception-oriented chart understanding tasks.
- Reasoning-Oriented Upgrades Yield Divergent Effects Across Chart Tasks. While DeepSeek-VL2 shows clear improvements over VL on Controlled Chart-to-Table, it exhibits a performance drop on Chart-to-Code, revealing a model-level trade-off introduced by enhanced reasoning capabilities. Controlled Chart-to-Table benefits from stronger high-level abstraction and structural planning, as it primarily involves localized perception and short-context reasoning without strict execution constraints. In contrast, Chart-to-Code requires faithful continuation of previously established object states. Under limited effective context length, VL2’s stronger reasoning tendency may reconstruct code structures based on inferred intent rather than preserved execution history, leading to logically plausible but non-executable outputs. This effect may be further amplified by VL2’s mixture-of-experts architecture, where dynamic expert routing can improve specialization for localized reasoning but introduces additional challenges in maintaining stable, long-horizon state consistency across generation steps. As a result, reasoning-centric and MoE-based upgrades can jointly benefit short-context understanding tasks while exacerbating failure modes in long-context, execution-sensitive generation.
@@
- Color Feature Presents Unique Challenges in Visual Decoding. Color accuracy remains the lowest-scoring aspect within the Visual Structure Consistency metrics, even for top-performing models such as GPT-5 (51.21), Claude-4-5-Sonnet(49.44) and Qwen3-VL-32B-Instruct(38.51). This suggests that fine-grained color differentiation poses unique challenges for current visual encoders, especially in complex or low-contrast chart regions. In many cases, the visual abstraction processes used by these models may reduce sensitivity to precise pixel-level color features.
@@
- High Pass Rate but Low Fidelity.Another failure mode is models generating syntactically valid outputs that “mask critical failures” in content. For example, in controlled chart-to-table task, DeepSeek-VL-7B successfully produces parsable code nearly every time ( about 98% pass rate) , yet its data extraction is almost nonexistent (strict F1 <1%). It often writes basic chart code that runs but does not capture the actual data or visual details. This strategy yields a high functional score but extremely low data fidelity, indicating the model is defaulting to trivial or placeholder outputs to avoid errors rather than truly understanding the chart.
@@
A High CLIP Score Does Not Guarantee Accurate Data Reconstruction. A model may generate a chart that is visually convincing—possessing the correct chart type, structure, and color distribution—yet fails to preserve the underlying data. For instance, Claude-4.5-Sonnet achieves a high CLIP score (78.93), yet its data fidelity (36.70) falls slightly below that of Qwen3-VL-235B (38.07), despite the latter having a significantly lower CLIP score (68.62). This divergence illustrates that visual similarity and data accuracy are distinct dimensions. As noted in the introduction, a model may "reproduce a chart’s appearance while silently altering the data." Consequently, CLIP scores and Data F1 metrics must be viewed as complementary: the former detects visual discrepancies, while the latter reveals semantic errors. Both are indispensable for a comprehensive evaluation of chart fidelity.
E.4 Analysis Across Chart Types
@@
Figure 8: Comparison of Overall Across Chart Types for Different Models on Chart-to-Code Task
@@
Figure 9: Comparison of Slight F1-Score Across Chart Types for Different Models on Controlled Chart-to-Table Task
@@
E.5 Correlation with Human Evaluation
To assess the extent to which our multi-level evaluation framework aligns with human judgments, we conduct a human preference study based on pairwise comparisons in the chart-to-code generation task, which reduces annotator calibration bias and enables more reliable assessment of relative quality differences than absolute scoring.
Data Sampling.
We randomly sample $N=300$ chart pairs from the ChartAnchor test set. For each pair $(C_{A},C_{B})$, two different models generate executable plotting code from the same ground-truth chart input, ensuring that any observed differences arise solely from model behavior rather than input variation.
Annotation Protocol.
We recruit domain experts with prior experience in data visualization and chart analysis as annotators. For each chart pair, annotators are presented with the reference chart image together with the two rendered candidate charts. Following the core dimensions of our evaluation framework, annotators independently indicate their preference (“A is better” or “B is better”) along three axes:
  • Visual Structure Consistency, focusing on chart type, layout, textual elements, and color usage;
  • Semantic Data Fidelity, emphasizing numerical accuracy and preservation of underlying data relationships;
  • Perceptual Quality, capturing overall visual plausibility and semantic coherence.
No ties are allowed, enforcing a strict preference decision for each dimension.
Agreement Metrics.
Rather than evaluating absolute quality, this study focuses on relative ranking alignment. For each chart pair and evaluation dimension, we compare the chart preferred by human annotators with the chart assigned a higher score by the corresponding automated metric. We report Consistency Accuracy (Acc) as the fraction of pairs for which the metric agrees with human preference. In addition, we compute Kendall’s rank correlation coefficient ($\tau$) to measure global agreement between metric-induced rankings and human preference orderings.
@@ Appendix F Case Study of ChartAnchor
Figure 10: Image Example1 for Chart-to-Code Task.
@@
Figure 11: Image Example2 for Chart-to-Code Task.
@@
Figure 12: Image Example3 for Chart-to-Code Task. The gray image indicates that the model did not generate a valid image or the code parsing failed
@@
Figure 13: Image Example4 for Chart-to-Code Task. The gray image indicates that the model did not generate a valid image or the code parsing failed
@@
Figure 14: Image Example5 for Chart-to-Code Task. The gray image indicates that the model did not generate a valid image or the code parsing failed
@@
Figure 15: Image Example6 for Chart-to-Code Task.
@@
Figure 16: Image Example7 for Chart-to-Code Task. Each subfigure shows the output from a different model.
@@
Figure 17: Image Example8 for Chart-to-Code Task. The gray image indicates that the model did not generate a valid image or the code parsing failed.
@@
Example 9: From the gold table (Table 11) and the generated tables (Figure 18), by comparing the row structures and numeric alignments, we observe that claude-3-7-Sonnet produces a table closely aligned with the gold table, with all values either correct or slightly off. Qwen2.5-VL-7B shows noticeable deviations in the "RDS 18–49" and "RDS Total" columns for 2014 and 2016. Qwen2.5-VL-32B gets closer but still introduces uniform values in 2015 that deviate from the gold data. Claude-3-7-Sonnet demonstrates the highest data fidelity and variation consistency across rows.
Table 11: Example 9: Gold Table
@@
Figure 18: Example9. Gold Image and Generated Tables for Controlled Chart-to-Table Task.
@@
Table 12: Example 10: Gold Table
Example 10: From the gold table (Table 12) and the generated tables (Figure 19), we can see, in this example, Claude-3-7-Sonnet is the only model that correctly identifies the categorical theta labels (e.g., AL29, GD35) and extracts reasonable numerical values for both "Today-r" and "Mean-r," achieving the highest alignment with the gold table. In contrast, Qwen2.5-VL-7B and Qwen2.5-VL-32B fail to capture the label names and instead generate numerical theta angles, indicating a misunderstanding of the scatterpolar chart structure. Their values also significantly diverge from the reference data, limiting their usability for structured analysis.
Figure 19: Example10. Gold Image and Generated Tables for Controlled Chart-to-Table Task.
Example 11: From the gold table (Table 13) and the generated tables (Figure 20), we can see, Claude-3-7-Sonnet is the only model that accurately captures both the structure and values of the gold table. It correctly identifies incremental changes (including positive and negative values) and matches the final "Price current" result. In contrast, Qwen2.5-VL-7B and Qwen2.5-VL-32B misinterpret the chart as cumulative values rather than stepwise deltas. As a result, their tables lose the core logic of a waterfall breakdown and deviate significantly from the reference.
@@
Table 13: Example 11: Gold Table
Figure 20: Example11. Gold Image and Generated Tables for Controlled Chart-to-Table Task.
@@ Appendix G Chart Examples
This section presents visual examples of the chart categories included in our benchmark. Each image illustrates a distinct chart type, showcasing the diversity in structure, data encoding, and visual design. Figures 21 and Figures 22 display representative samples from all 30 categories.
Figure 21: Representative chart images in our benchmark.
Figure 22: Representative chart images in our benchmark.
Appendix H Volunteer Recruitment and Payment
All annotations were completed by five student volunteers from our institution. Given the light workload of the annotation task, participation was voluntary and uncompensated.
Appendix I Data Consent
All data used in this work are publicly accessible and freely downloadable from their respective websites.
Appendix J AI Assistance Disclosure
AI assistants were used solely for editorial assistance, including grammar and wording refinement, with all substantive content authored and reviewed by the authors.
Appendix K Data Screening and Privacy Protection
We verified that all datasets used in this work are publicly available and do not contain personally identifiable information or offensive content. As the data are sourced from curated public repositories and websites, no additional anonymization or filtering was required.